# Ernesto L. Estevanell-Valladares

### v1

@@ 5 Discussion
Warm-start priors consistently steer the search toward stronger performance–time trade-offs across all six benchmarks. Figure 1 reports the winning ratio: the share of evaluated LM pipelines that improve upon the zero-shot Pareto front. The High–Max prior is the most stable, winning about 20% of pipelines on SQuAD, LIAR, MELD, and DROP, and 10–15% on SST-2 and AG News. On the LIAR and MELD pair, the High–LIAR prior achieves winning ratios near 50% and 40%, respectively, while cutting the error rate by a factor of seven (Table 5).
Figure 1: Ratio of discovered LM pipelines outperforming the Zero-shot baseline in Text Classification and QA.
@@
Figure 3: Distance between Text Classification Tasks according to their meta-features (§3.2.1).
These results show that combining experience discrimination with adaptive probability shifts yields the best of both worlds: rapid convergence when relevant meta-knowledge exists yet robustness when it does not. Whenever the experience store contained closely related traces—e.g., MELD–LIAR (Figure 3)—the similarity-aware priors trimmed average evaluation time by up to 4.5x and increased peak $F_{1_{\text{m}}}$ (Table 5). Even on sparsely related tasks such as SST-2 and AG News, softer pulls uncovered superior Pareto trade-offs by moderating exploration strength (Figure 2(a)).
The baseline performance of kNN highlights the significance of selective memory. While it has access to both positive and negative examples, it assigns equal weight to all neighbours, failing to demote weak configurations and causing accuracy to fall on three of four classification datasets. In contrast, XAutoLM’s asymmetric pull–push update penalises both past failures and underperforming successes. DROP, for example, illustrates the need to learn from failures: a low-bias prior that ignores negatives collapses to $F_{1}=0.18$, whereas reinstating the push restores $F_{1}=0.40$ and halves mean evaluation time.
Our findings further show that transfer using our method extends beyond classification. With barely a handful of relevant experience, a high-bias prior multiplies SQuAD $F1$ from ~0.3 to nearly 0.9 and compresses evaluation time by threefold, producing a dominant Pareto front (Figure 2(b)). On the other hand, DROP illustrates the importance of negative experiences: a low-bias prior that ignores negatives collapses to $F_{1}=0.18$, whereas reinstating the push restores $F_{1}=0.40$ and cuts mean evaluation time by 50 % (Table 6).
A core motivation of our framework is to reduce the carbon footprint and environmental toll of repeated large-scale language model fine-tuning. By systematically reusing insights from past runs, XAutoLM significantly reduces redundant evaluations and lowers the overall error rate during the search. Beyond simply lowering compute hours, this approach aligns with the growing Green AI ethos in NLP (Wang et al., 2023b; Schwartz et al., 2020), emphasising the importance of responsible resource usage. Our experiments illustrate that our warm-start strategy improves performance and streamlines the search process, yielding algorithms that better balance efficiency and performance.
@@ 6 Conclusions
XAutoLM converts the costly trial–and–error of language model fine-tuning into a guided, resource-aware search. By seeding the optimiser with a similarity-weighted prior built from past successes & failures, the framework consistently uncovers superior performance–time trade-offs. Across four text-classification corpora and two generative QA benchmarks, it surpasses the best zero-shot $F_{1}$ on five tasks—matching it on SST-2—while cutting mean evaluation time by up to a factor of four and reducing error rates by as much as sevenfold. These gains hold across a refreshed model pool that ranges from lightweight discriminative to compact, state-of-the-art compact generation models. Because every recovered pipeline reuses information already paid for, XAutoLM advances the Green AI agenda (Schwartz et al., 2020), delivering competitive results while avoiding redundant computation.
@@ 7 Limitations
[…]
@@ A.3 Beta Scale and Utility Functions
All candidates for the single-objective experiments (Appendix C) utilise a weighted sum approach with the $F1$ score weight set to 1 and the evaluation time weight set to 0. Detailed specifications of candidate configurations can be found in the visualizations provided in the respective sections (Appendix C for single-objective, and Appendix B for multi-objective).
@@ A.4 Experimental Setup and Computational Resources
The main text fully discloses our experimental setup (§4).
@@ A.5 Framework Overview and Dependencies
XAutoLM is implemented on top of the AutoGOAL framework Estevanell-Valladares et al. (2024); Estevez-Velarde et al. (2020), leveraging its optimisation strategy and abstractions. Our implementation is developed in Python and utilizes the HuggingFace Transformers library Wolf et al. (2019) to access pre-trained language models. A complete list of dependencies, environment setup instructions, and detailed documentation on how to run the experiments (and statistical testing), reproduce the results, and navigate the codebase is provided in the repository.
The code and all associated materials can be accessed at the following anonymous GitHub repository: https://anonymous.4open.science/r/XAutoLLM-A010 (currently private for blind review; will be made public upon completion of the review process).
@@ Appendix B Multi-Objective Initial Probabilities
This appendix visualises the initial probability distributions over fine-tuning methods induced by different meta-learning configurations (Prior) in our multi-objective experiments (see §4). Each configuration is defined by:
@@
    $\alpha_{\max}^{+}$ and $\alpha_{\max}^{-}$ values (fixed or adaptive) (§3.3).
Each configuration yields a distinct initial probability vector for the available fine-tuning methods, with deviations from the baseline distribution measured via Total Variation (TV). Grouping configurations by TV allows us to categorise them into low, moderate, and high bias levels relative to the baseline’s uniform initialisation.
@@ B.1 Classification Tasks
Figure 6 shows the meld dataset’s initial distributions. As discussed in § 4, meld shares some meta-feature similarities with liar (Figure 3), causing some distributions to concentrate around methods found promising in liar’s prior runs.
@@ B.2 QA Tasks
These visualisations underscore how our meta-learning strategy adapts the search space before optimisation begins. By systematically adjusting the initial probabilities, XAutoLM avoids mindlessly searching all possibilities and exploits task similarities to emphasise configurations that are historically more successful or resource-feasible.
@@ Appendix C Single-Objective Warm Start Evaluation
[…]
@@ C.2 Performance Results
Table 7 reports our results. We conducted a detailed statistical analysis across six independent runs per configuration on LIAR and SST-2, evaluating performance, convergence time, and reliability. Normality was tested using Shapiro–Wilk, followed by ANOVA McHugh (2011) for normal metrics, and Friedman tests (Pereira et al., 2015) for non-parametric ones. We report Cohen’s $d$ and Cliff’s $\delta$ as effect-size measures; power analyses accompany each test in the repository.
@@
On LIAR, while none of the warm-start priors significantly outperformed the baseline in peak $F1_{\text{macro}}$ (ANOVA $p=0.856$, Friedman $p=0.94$), we observed a significant overall improvement in mean performance across groups (ANOVA $p=0.005$, Friedman $p=0.004$). Post-hoc comparisons, however, were not significant after correction, likely due to limited sample size. More notably, the error ratio—the share of failed evaluations—dropped dramatically from 0.69 (baseline) to 0.24 (High WS), a difference found to be statistically significant (Friedman $p=0.031$) with a large effect size (Cohen’s $d=3.39$). Convergence time metrics (TT50, TT75, TT90) also trended lower, with moderate effect sizes, although these differences did not reach statistical significance.
@@
In summary, warm-start priors consistently yielded practical convergence speed and robustness benefits. While not all improvements were statistically significant—expected under a small-sample regime—our analysis shows that key metrics such as error ratio and mean F1 on LIAR and max F1 on SST-2 do reach significance. Full results, post hoc comparisons, and power analyses are available in our open-source repository.
@@ Appendix D Pareto Front Visualizations
[…]
@@
[…]

### v2

@@ 5 Discussion
Warm-start priors consistently steer the search toward stronger performance–time trade-offs across all six benchmarks. Figure 1 reports the winning ratio: the share of evaluated LM pipelines that improve upon the zero-shot Pareto front.
Figure 1: Ratio of discovered pipelines outperforming the Zero-shot baseline in Text Classification and QA.
@@
The High–Max prior is the most stable, winning about 20% of pipelines on SQuAD, LIAR, MELD, and DROP, and 10–15% on SST-2 and AG News. On the LIAR and MELD pair, the High–LIAR prior achieves winning ratios near 50% and 40%, respectively, while cutting the error rate by a factor of seven (Table 5). For clarity, all ET values are computed only on successful evaluations, while failure rates are captured by the Error Ratio, with all methods allotted an identical 24 GPU‑hour wall‑clock budget per run.
Figure 3: Distance between Text Classification Tasks according to their meta-features (Section 3.2.1).
These results show that combining experience discrimination with adaptive probability shifts yields the best of both worlds: rapid convergence when relevant meta-knowledge exists yet robustness when it does not. Whenever the experience store contained closely related traces, e.g., MELD–LIAR (Figure 3), the similarity-aware priors trimmed average evaluation time by up to 4.5x and increased peak $F_{1_{\text{m}}}$ (Table 5). Even on sparsely related tasks such as SST-2 and AG News, softer pulls uncovered superior Pareto trade-offs by moderating exploration strength (Figure 2(a)).
The baseline performance of kNN highlights the significance of selective memory. While it has access to both positive and negative examples, it assigns equal weight to all neighbors, failing to demote weak configurations and causing accuracy to fall on three of four classification datasets. In contrast, XAutoLM’s asymmetric pull–push update penalizes both past failures and underperforming successes. DROP, for example, illustrates the need to learn from failures: a low-bias prior that ignores negatives collapses to $F_{1}=0.18$, whereas reinstating the push restores $F_{1}=0.40$ and halves mean evaluation time.
Our findings further show that transfer using our method extends beyond classification. With barely a handful of relevant experience, a high-bias prior multiplies SQuAD $F1$ from $\approx 0.3$ to $\approx 0.9$ and compresses evaluation time by threefold, producing a dominant Pareto front (Figure 2(b)). On the other hand, DROP illustrates the importance of negative experiences: a low-bias prior that ignores negatives collapses to $F_{1}=0.18$, whereas reinstating the push restores $F_{1}=0.40$ and cuts mean evaluation time by 50 % (Table 6).
A core motivation of our framework is to reduce the carbon footprint and environmental toll of repeated large-scale language model fine-tuning. By systematically reusing insights from past runs, XAutoLM significantly reduces redundant evaluations and lowers the overall error rate during the search. Beyond simply lowering compute hours, this approach aligns with the growing Green AI ethos in NLP (Wang et al., 2023b; Schwartz et al., 2020), emphasizing the importance of responsible resource usage. Our experiments demonstrate that our warm-start strategy enhances performance and streamlines the search process, resulting in algorithms that strike a better balance between efficiency and performance.
@@ 6 Conclusions
XAutoLM converts the costly trial–and–error of language model fine-tuning into a guided, resource-aware search. By seeding the optimizer with a similarity-weighted prior built from past successes & failures, the framework consistently uncovers pipelines with superior performance–time trade-offs. Across four text-classification corpora and two generative QA benchmarks, it surpasses the best zero-shot $F_{1}$ on five tasks, matching it on SST-2, while cutting mean pipeline evaluation time by up to a factor of four and reducing error rates by as much as sevenfold. These gains hold across a refreshed model pool that ranges from lightweight discriminative to compact generative models. Because every recovered pipeline reuses information already paid for, XAutoLM advances the Green AI agenda (Schwartz et al., 2020), delivering competitive results in less search time, while avoiding redundant computation.
@@ 7 Limitations
Scaling to bigger LLMs
XAutoLM is scale-agnostic: the optimizer treats candidates as black-box fit/evaluate calls and does not rely on model internals. Our open-source implementation presently evaluates on a single GPU, which constrained the largest models tested; this is a property of the evaluator backend, not of the optimization method. The experience store logs a minimal hardware profile (Section 3.2), which helps steer the search away from infeasible pipelines under a given machine with a single GPU setup. Supporting larger models, therefore, amounts to adding multi-GPU meta-features and swapping in a larger-model evaluator (e.g., parameter-efficient Hu et al. (2021)/quantized Nagel et al. (2021); Dettmers et al. (2023) or distributed evaluators Zhao et al. (2023)) in future releases; the search algorithm and experience-based priors remain unchanged. We leave such engineering backends to future work and keep our claims limited to the single-GPU setting evaluated here.
Multimodality
The current experience store and benchmarks are text-only; verifying that the warm-start prior transfers to dialogue, speech, or multimodal pipelines is an essential next step.
Statistical Tests
Statistical support is available only for the single-objective probes archived in Appendix C. Extending significance testing to the multi-objective fronts of Tables 5 and 6 would require many repeated runs and is left for future work, where bootstrap or fully Bayesian analyses are planned.
Efficiency Measures
Our energy discussion rests on the empirical link between execution time and power draw reported by prior work Wang et al. (2023b); Estevanell-Valladares et al. (2024); we did not log wattage directly. The next release of XAutoLM will record real-time power and emit CO2 estimates alongside performance metrics.
Acknowledgments
This research has been partially funded by the University of Alicante, the University of Havana, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: At the regional level, and as the primary source of support, the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), FEDER granted funding for CIDEGENT (CIDEXG/2023/13); and NL4DISMIS (CIPROM/2021/21). At the national level, the following projects were granted: HEART-NLP (PID2024-156263OB-C22); COOLANG (PID2021-122263OB-C22); SOCIALTRUST (PDC2022-133146-C22); ILENIA (2022/TL22/00215334) and ALIA models (https://alia.gob.es) funded by MCIN/AEI/10.13039/501100011033 and, as appropriate, by ERDF A way of making Europe, by the European Union or by the European Union NextGenerationEU/PRTR; and by the State Subprogram for Training, Attraction, and Retention of Talent (PEICTI 2024) of the Spanish Ministry of Science and Innovation, grant PRX24/00272.
@@ A.3 Beta Scale and Utility Functions
All candidates for the single-objective experiments (Appendix C) utilize a weighted sum approach with the $F1$ score weight set to 1 and the evaluation time weight set to 0. Detailed specifications of candidate configurations can be found in the visualizations provided in the respective sections (Appendix C for single-objective, and Appendix B for multi-objective).
@@ A.4 Experimental Setup and Computational Resources
The main text fully discloses our experimental setup (Section 4).
@@ A.5 Framework Overview and Dependencies
XAutoLM is implemented on top of the AutoGOAL framework Estevanell-Valladares et al. (2024); Estevez-Velarde et al. (2020), leveraging its optimization strategy and abstractions. Our implementation is developed in Python and utilizes the HuggingFace Transformers library Wolf et al. (2019) to access pre-trained language models. A complete list of dependencies, environment setup instructions, and detailed documentation on how to run the experiments (and statistical testing), reproduce the results, and navigate the codebase is provided in the repository.
The code and all associated materials can be accessed at the following GitHub repository: https://github.com/EEstevanell/XAutoLM.
@@ Appendix B Multi-Objective Initial Probabilities
This appendix visualizes the initial probability distributions over fine-tuning methods induced by different meta-learning configurations (Prior) in our multi-objective experiments (see Section4). Each configuration is defined by:
@@
    $\alpha_{\max}^{+}$ and $\alpha_{\max}^{-}$ values (fixed or adaptive) (Section 3.3).
Each configuration yields a distinct initial probability vector for the available fine-tuning methods, with deviations from the baseline distribution measured via Total Variation (TV). Grouping configurations by TV allows us to categorize them into low, moderate, and high bias levels relative to the baseline’s uniform initialisation.
@@ B.1 Classification Tasks
Figure 6 shows the meld dataset’s initial distributions. As discussed in Section 4, meld shares some meta-feature similarities with liar (Figure 3), causing some distributions to concentrate around methods found promising in liar’s prior runs.
@@ B.2 QA Tasks
These visualizations underscore how our meta-learning strategy adapts the search space before optimization begins. By systematically adjusting the initial probabilities, XAutoLM avoids mindlessly searching all possibilities and exploits task similarities to emphasize configurations that are historically more successful or resource-feasible.
@@ Appendix C Single-Objective Warm Start Evaluation
This appendix reports single-objective experiments optimizing the macro-$F1$ score alone. We compare the Zero-shot AutoGOAL baseline against three representative warm-start priors, Low, Moderate, and High bias, selected from fourteen candidate configurations grouped by total variation (TV) distance. All priors use the std-only $\beta$ scale, Euclidean distance, and fixed learning rates ($\alpha_{\max}^{+}=0.05$, $\alpha_{\max}^{-}=0.02$).
@@ C.2 Performance Results
Table 7 reports our results. We conducted a detailed statistical analysis across six independent runs per configuration on LIAR and SST-2, evaluating performance, convergence time, and reliability. Normality was tested using Shapiro–Wilk, followed by ANOVA McHugh (2011) for normal metrics, and Friedman tests (Pereira et al., 2015) for non-parametric ones. We report Cohen’s $d$ and Cliff’s $\delta$ as effect‐size measures; power analyses accompany each test in the repository.
@@
On LIAR, while none of the warm-start priors significantly outperformed the baseline in peak $F1_{\text{macro}}$ (ANOVA $p=0.856$, Friedman $p=0.94$), we observed a significant overall improvement in mean performance across groups (ANOVA $p=0.005$, Friedman $p=0.004$). Post-hoc comparisons, however, were not significant after correction, likely due to limited sample size. More notably, the error ratio, the share of failed evaluations, dropped dramatically from 0.69 (baseline) to 0.24 (High WS), a difference found to be statistically significant (Friedman $p=0.031$) with a large effect size (Cohen’s $d=3.39$). Convergence time metrics (TT50, TT75, TT90) also trended lower, with moderate effect sizes, although these differences did not reach statistical significance.
@@
In summary, warm-start priors consistently yielded practical convergence speed and robustness benefits. While not all improvements were statistically significant, expected under a small-sample regime, our analysis shows that key metrics such as error ratio and mean F1 on LIAR and max F1 on SST-2 do reach significance. Full results, post hoc comparisons, and power analyses are available in our open-source repository.
@@ Appendix D Pareto Front Visualizations
Figure 10: Comparison of Pareto fronts for zero-shot baseline (solid blue line) and warm-start priors at Low (green), Moderate (orange), and High (red) bias levels. Each point plots $(ET,F1_{\text{macro}})$ for classification tasks (a–d) or $(ET,F1)$ for QA tasks (e–f). Points to the left or above the baseline outperforms the zero-shot Pareto front.
@@
High-bias priors calibrated on LIAR produce up to 40% of pipelines that dominate the baseline, reducing error rates by roughly sevenfold (cf. Table 5). Due to the substantial meta-feature similarity between LIAR and MELD (Figure 3), both tasks see rapid convergence to high-$F1_{\text{macro}}$ regions.